OpenStudy (anonymous):

Eliminate the “z” variable in the first two equations by finding opposite coefficients and adding. What is the resulting equation? 4x – y + z = 6 3x + 2y – z = 0 4x – y – 3z = 2 7x + y = 6 x – y = 6 x + y = 6 7x – y = 6

OpenStudy (kinggeorge):

In the first two equations, the \(z\)'s already have opposite coefficients. So you just need to add the two equation.

OpenStudy (kinggeorge):

So can you find \[\begin{align} &4x-y+z=6 \\ + \qquad &3x+2y-z=0\\&\text{______________}\end{align}\]